In an isobaric process the ratio of heat supplied to the system (dq) and workdone by the system (dw) for diatomic gas is

Grade:12th pass

3 Answers

Vikas TU
7 years ago
For the isobaric process, for n = 1 mol.
From the first law of thermodynamics,
U = 2.5*RT
W = R*T
and Q => U – W => 2.5RT – RT => 0.5RT J
Khimraj
7 years ago
For isobaric process and n = 1. dQ = 5/2 RdT + RdT = 7/2RdT. And dW = RdT. So ratio of dQ/dT is 7/2.Hope it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.
Khimraj
7 years ago
Sorry for correction as in above solution i assume question of physics. For chemistry sol is as.For isobaric process and n = 1. dQ = 5/2 RdT - RdT = 3/2RdT. And dW = RdT. So ratio of dQ/dT is 3/2.Hope now it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.